OpenStudy (anonymous):

REWARD--->Single die is rolled whats the probability of rolling an even # or a number less than 4!?!?!?:/

OpenStudy (anonymous):

P(even no.)= 3/6 P(no<4)= 3/6 3/6+3/6=1

OpenStudy (anonymous):

wrong it was 2/5 thanks tho